Prove that n + log2n = O(n) by showing that there exists a constant c > 0 such that n + log2n ≤ cn.
(note that log2n means (log n)2.)
1
Expert's answer
2019-02-22T05:38:07-0500
"(\\log^2\u2061n-n)'=2 \\log \u2061n \\times 1\/n-1=2 ((\\log \u2061n-n))\/n\\le0,\\quad n\\in\\mathbb{N}""(\\log \u2061n-n)'=1\/n-1\\le0""\log^2 n \le n""n + \log^2 n \le n + n = 2n,\quad n\in \mathbb{N}"